Circular No. 8016

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hy IN NGC 3464 L. A. G. Monard, Pretoria, South Africa, reports the discovery of an

apparent supernova (mag 16.4 ± 0.3) on unfiltered CCD images taken on Nov. 12.098 UT with a 0.30-m reflector; it was confirmed at the same brightness on an image taken on Nov. 14.084. SN 2002hy is located at α = 10^{h}54^{m}39^{s}.18, δ = −21^{o}03′41′′.2 (equinox 2000.0), which is 14′′.8 west and 17′′.0 north of the nucleus of NGC 3464. Nothing is visible at this location on the Digitized Sky Survey (limiting red mag 19.5) or on an image taken by Monard on 2002 Oct. 13.104 (limiting mag 18.0). SN 2002J (cf. IAUC 7800) also appeared in NGC 3464.

SUPERNOVAE 2002gf, 2002gg, AND 2002hx A. V. Filippenko and R. Chornock, University of California, Berkeley,

report that inspection of CCD spectra (range 310–1000 nm) obtained on Nov. 11 UT with the Keck I 10-m telescope (+ LRIS) shows that SN 2002gf and SN 2002gg (IAUC 7990) are of type Ia, ∼ 1–1.5 months past maximum brightness. The redshifts of the host galaxies, measured from narrow emission lines (H II regions) in the nuclei, are 0.086 and 0.110, respectively.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002hx (cf. IAUC 8015), obtained by M. Calkins on Nov. 14.53 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. The spectrum consists of a blue continuum with P-Cyg lines of hydrogen. Adopting the NASA/IPAC Extragalactic Database recession velocity of 9293 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 8800 km/s.

V838 MONOCEROTIS T. R. Geballe, Gemini Observatory; and B. Smalley, A. Evans, and

M. T. Rushton, Keele University, report on spectra obtained at the U.K. Infrared Telescope (+ CGS4) on Oct. 29.125 UT, at resolutions of 0.0015 and 0.0030 μm over ranges 0.80–1.34 and 1.45–2.52 μm, respectively: “Extraordinary changes have occurred since our last reported measurements (IAUC 7796). The spectrum is now that of an extremely cool supergiant, with very deep water bands (reminiscent of those seen in T-type brown dwarfs) and remarkably strong CO first-overtone bands. There are additional strong molecular bands present at 0.85, 0.89, 0.97, 1.05, 1.23, 1.25, 1.65, and 1.69 μm. Further observations are urged.”

2002 November 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green